# Exploring Human-in-the-Loop Test-Time Adaptation by Synergizing Active Learning and Model Selection

### v2

@@ Abstract
Existing test-time adaptation (TTA) approaches often adapt models with the unlabeled testing data stream. A recent attempt relaxed the assumption by introducing limited human annotation, referred to as Human-In-the-Loop Test-Time Adaptation (HILTTA) in this study. The focus of existing HILTTA studies lies in selecting the most informative samples to label, a.k.a. active learning. In this work, we are motivated by a pitfall of TTA, i.e. sensitivity to hyper-parameters, and propose to approach HILTTA by synergizing active learning and model selection. Specifically, we first select samples for human annotation (active learning) and then use the labeled data to select optimal hyper-parameters (model selection). To prevent the model selection process from overfitting to local distributions, multiple regularization techniques are employed to complement the validation objective. A sample selection strategy is further tailored by considering the balance between active learning and model selection purposes. We demonstrate on 5 TTA datasets that the proposed HILTTA approach is compatible with off-the-shelf TTA methods and such combinations substantially outperform the state-of-the-art HILTTA methods. Importantly, our proposed method can always prevent choosing the worst hyper-parameters on all off-the-shelf TTA methods. The source code will be released upon publication.
**footnotetext: Equal contribution.footnotetext: Corresponding author. Email: alex.xun.xu@gmail.com
@@ 1 Introduction
Despite the theoretical guarantee and superior empirical observations, the underlying practice of utilizing label budget by existing active HILTTA methods follows the traditional line of research on active learning (Sener & Savarese, 2018; Saran et al., 2023). In this study, we aim to explore the question: How can human-in-the-loop test-time adaptation be realized beyond an active learning perspective? To answer this question, we first revisit the pitfalls of state-of-the-art test-time adaptation (Zhao et al., 2023). Among the multiple remaining pitfalls, we pinpoint hyper-parameter selection as the most nasty one due to the sensitivity to the choice of hyper-parameters by many TTA methods. Hyper-parameter selection, also referred to as model selection, aims to choose the optimal hyper-parameters for model training. Established paradigms towards model selection often rely on a held-out validation set and the model selection task is defined as discovering the hyper-parameters trained upon which the model displays optimal generalization performance on the held-out validation set (Yang & Shami, 2020). Inspired by the success of hyper-parameter optimization, we offer a fresh perspective on human-in-the-loop test-time adaptation by synergizing active learning and model selection. In specific, we present an HILTTA procedure with three steps. First, upon observing a testing sample, the model will make an instant prediction to minimize inference latency. When a fixed number of testing samples are observed, we employ a data selection strategy to recommend a subset of testing samples for human oracle to label. The labeled data will then be utilized as the validation set for selecting the best hyper-parameters. Finally, the model is further updated with the labeled data in a fully supervised fashion. With the novel HILTTA steps, we intend to make better utilization of the limited annotation budget compared to existing HILTTA approaches.
@@ 2.2 Active Learning
Active learning (AL) aims to select the most informative samples for Oracle to label for supervised training. Prevailing approaches often prioritize uncertainty (Gal & Kendall, 2017), feature diversity (Sener & Savarese, 2018) or the combination of both (Ash et al., 2020). The traditional AL approaches often adopt a batch-wise paradigm. When data arrive in a sequential manner stream-based AL (Saran et al., 2023) proposed to quantify the increment of the determinant as the criteria for incremental sample selection. Active learning was integrated into test-time adaptation to tackle the forgetting issue (Gui et al., 2024). In contrast to tackling HILTTA from a pure active learning perspective, we propose to synergize active learning and model selection. Selecting samples for model selection often relies on an orthogonal objective to active learning. In the related works, active testing (Kossen et al., 2021; 2022) was developed to select a subset of testing data that is representative of the whole testing data. Active testing has been demonstrated to help label-efficient model validation. Nevertheless, active learning and active testing will prioritize supervised training and model validation respectively. We believe either is not enough for the purpose of synergizing active learning and model selection for test-time adaptation.
@@ 3.4 Sample Selection for HILTTA
The sample selection strategy should balance the objectives of active learning and model selection for more effective HILTTA. We achieve this target from two perspectives. First, the selected samples should prioritize low-confidence samples for more effective active learning, as adopted by SimATTA (Gui et al., 2024) through incremental K-means. However, selecting redundant low-confidence samples may harm the effectiveness of model validation. This motivates us to select samples approximating the testing data distribution, similar to the objective of active testing (Kossen et al., 2021). Biasing towards either criterion would result in a suboptimal HILTTA.
The above objectives could be achieved by a recent active learning approach, BADGE (Ash et al., 2020), which takes the gradient of loss w.r.t. the penultimate layer feature as the uncertainty weighted feature as $g_{x}=\frac{\partial}{\partial\theta_{\theta_{out}}}\mathcal{L}_{ce}(h(x;\theta),\hat{y})$, on which K-Means++ is applied for active selection. The feature dimension is the product of the number of classes $C$ and penultimate layer channels $D$, i.e. $g_{x}\in\mathbbm{R}^{C\times D}$. Accumulating the feature overall categories results in overly high dimension features for large-scale classification tasks, e.g. $C=1000$ for ImageNet (Deng et al., 2009). To mitigate the ultra-high dimension challenge, we integrate uncertainty into feature representation by multiplying the uncertainty, measured by the margin confidence, to the feature representation as in Eq. 7, where $\mathrm{sort}^{d}$ refers to the descending sorting.
@@
The above design prioritizes the difference between the probability of the most confident and second most confident classes as a quantification of model uncertainty.
@@ 4.1 Experimental setting
We use the optimizer recommended in the original papers for each method’s unsupervised training. For supervised training, we use the Adam optimizer with a learning rate of 1e-5. We consider seven distinct values for each hyper-parameter category regarding model selection, detailed in Tab. 6.
@@ 4.2 Evaluations on HILTTA
Table 1: Average classification error (lower is better) through ongoing adaptation to varying domains. "Worst", "Avg.", and "Best" indicate the worst, average, and best performance within the candidate hyper-parameter pool $\Omega$. The improvement of error rate for each off-the-shelf TTA method after combining with HIL is appended in the bracket as ($\delta$). * denotes methods under manually tuned hyper-parameters. $\diamondsuit$ denotes methods combined with our proposed HILTTA (with model selection).
@@ 4.3 Ablation & Additional Study
Evaluations beyond 2D Image Recognition: To further evaluate the effectiveness of our method beyond 2D image classification task, we conduct experiments on 3D point cloud classification using DGCNN (Wang et al., 2019) as the backbone, adapting it continuously across 15 domains within the ModelNet40-C (Sun et al., 2022) dataset. We maintained the same hyper-parameter candidate set as outlined in Table 6 and followed the experimental setup detailed in Table 1, with a batch size of 32 and an annotation percentage of 3.2%. The results, presented in Figure 3, show that unsupervised TTA still consistently outperforms the baseline without any adaptation (Source). However, due to the sensitivity to hyper-parameters, all unsupervised TTA methods exhibit a substantial variation in performance with the “Best” and “Worst” hyper-parameters. All off-the-shelf TTA methods are consistently improved when integrated with our HILTTA method. Five out of the six methods witness a more significant boost in performance, surpassing their unsupervised counterpart with “Best” hyper-parameters.
@@ Appendix B Additional Experimental Analysis
Studies for Regularization in HILTTA: We evaluate the impact of the proposed anchor regularization and EMA smoothing techniques for model selection within the HILTTA framework. These methods are designed to prevent the model from overfitting to the local data distribution in the testing data stream. While removing these regularizations may improve adaptation to a single domain, it negatively impacts performance across future domains. As shown in Tab. 7, in the continual TTA setting, model selection without regularization results in significantly worse performance compared to when regularization is applied (17.25% vs. 15.87%). This indicates that regularization is crucial for maintaining robust performance in continual TTA. , model selection without regularization leads to significantly worse results (17.25% vs. 15.87%), highlighting the necessity of regularization for robust performance in continual TTA.
@@
Comparing with Knowledge Distillation and Ensemble Learning: We carried out additional experiments by comparing with ensemble learning and multi-teacher knowledge distillation methods. Specifically, for ensemble learning, we average the posterior predicted by each candidate model and make a prediction with the average posterior. For multi-teacher knowledge distillation, we use the pseudo label predicted by the ensemble model for self-training. As observed from Tab. 8, both ensemble learning and multi-teacher knowledge distillation yields inferior results than HILTTA. This is most likely caused by incorporating poor candidate models (ensemble learning) and using poor pseudo labels for self-training, i.e. learning from noisy labeled data (multi-teacher knowledge distillation).
@@
Comparing Adaptation Time & Annotation Time: We carried out empirical studies by measuring the wall-clock time lapses of the inference step (Inference Time), adaptation step (Adaptation Time), and human annotation (Annotation Time) per sample. In specific, the "Inference Time" refers to the time-lapse of pure forward pass, the "Adaption Time" refers to the time-lapse of backpropagation and weights update, and "Annotation Time" refers to the estimated time lapse required for human to annotate images at a 3% labeling rate. We evaluate under three categories of annotation cost, i.e. 2 sec/sample, 5 sec/sample, and 10 sec/sample. As seen from Tab. 9, despite the adaptation time being higher than the inference time (8.7 ms v.s. 0.4 ms), the annotation time is the most expensive step. These results suggest model selection and adaptation do not slow down the throughput under the HILTTA protocol.

### v3

@@ Abstract
Existing test-time adaptation (TTA) approaches often adapt models with the unlabeled testing data stream. A recent attempt relaxed the assumption by introducing limited human annotation, referred to as Human-In-the-Loop Test-Time Adaptation (HILTTA) in this study. The focus of existing HILTTA studies lies in selecting the most informative samples to label, a.k.a. active learning. In this work, we are motivated by a pitfall of TTA, i.e. sensitivity to hyper-parameters, and propose to approach HILTTA by synergizing active learning and model selection. Specifically, we first select samples for human annotation (active learning) and then use the labeled data to select optimal hyper-parameters (model selection). To prevent the model selection process from overfitting to local distributions, multiple regularization techniques are employed to complement the validation objective. A sample selection strategy is further tailored by considering the balance between active learning and model selection purposes. We demonstrate on 5 TTA datasets that the proposed HILTTA approach is compatible with off-the-shelf TTA methods and such combinations substantially outperform the state-of-the-art HILTTA methods. Importantly, our proposed method can always prevent choosing the worst hyper-parameters on all off-the-shelf TTA methods. The source code is available at https://github.com/Yushu-Li/HILTTA.
**footnotetext: Equal contribution. This work was done during Yushu Li and Yongyi Su’s attachment with $I^{2}R$ as visiting students.footnotetext: Corresponding author. Email: alex.xun.xu@gmail.com
@@ 1 Introduction
Table 1: The differences between our proposed human-in-the-loop test-time adaptation (HILTTA) and existing TTA settings.
Setting Active Learning Model Selection
Test-Time Adaptation (TTA) - -
Active Test Time Adaptation (ATTA) $\checkmark$ -
Human-in-the-Loop TTA (HILTTA) $\checkmark$ $\checkmark$
Despite the theoretical guarantee and superior empirical observations, the underlying practice of utilizing label budget by existing active HILTTA methods follows the traditional line of research on active learning (Sener & Savarese, 2018; Saran et al., 2023). In this study, we aim to explore the question: How can human-in-the-loop test-time adaptation be realized beyond an active learning perspective? To answer this question, we first revisit the pitfalls of state-of-the-art test-time adaptation (Zhao et al., 2023). Among the multiple remaining pitfalls, we pinpoint hyper-parameter selection as the most nasty one due to the sensitivity to the choice of hyper-parameters by many TTA methods. Hyper-parameter selection, also referred to as model selection, aims to choose the optimal hyper-parameters for model training. Established paradigms towards model selection often rely on a held-out validation set and the model selection task is defined as discovering the hyper-parameters trained upon which the model displays optimal generalization performance on the held-out validation set (Yang & Shami, 2020). Inspired by the success of hyper-parameter optimization, we offer a fresh perspective on human-in-the-loop test-time adaptation by synergizing active learning and model selection. In specific, we present an HILTTA procedure with three steps. First, upon observing a testing sample, the model will make an instant prediction to minimize inference latency. When a fixed number of testing samples are observed, we employ a data selection strategy to recommend a subset of testing samples for human oracle to label. The labeled data will then be utilized as the validation set for selecting the best hyper-parameters. Finally, the model is further updated with the labeled data in a fully supervised fashion. With the novel HILTTA steps, we intend to make better utilization of the limited annotation budget compared to existing HILTTA approaches. We demonstrate the differences between our proposed HILTTA and related TTA settings in Tab. 1.
@@ 2.2 Active Learning
Active learning (AL) aims to select the most informative samples for Oracle to label for supervised training. Prevailing approaches often prioritize uncertainty (Gal & Kendall, 2017), feature diversity (Sener & Savarese, 2018) or the combination of both (Ash et al., 2020). The traditional AL approaches often adopt a batch-wise paradigm. When data arrive in a sequential manner stream-based AL (Saran et al., 2023) proposed to quantify the increment of the determinant as the criteria for incremental sample selection. Active learning was integrated into test-time adaptation to tackle the forgetting issue (Gui et al., 2024). In contrast to tackling HILTTA from a pure active learning perspective, we propose to synergize active learning and model selection. Selecting samples for model selection often relies on an orthogonal objective to active learning. In the related works, active testing (Kossen et al., 2021; 2022; Matsuura & Hara, 2023; Ochiai et al., 2023) was developed to select a subset of testing data that is representative of the whole testing data. Active testing has been demonstrated to help label-efficient model validation. Nevertheless, active learning and active testing will prioritize supervised training and model validation respectively. We believe either is not enough for the purpose of synergizing active learning and model selection for test-time adaptation.
@@ 3.4 Sample Selection for HILTTA
The sample selection strategy should balance the objectives of active learning and model selection for more effective HILTTA. We achieve this target from two perspectives. First, the selected samples should prioritize low-confidence samples for more effective active learning, as adopted by SimATTA (Gui et al., 2024) through incremental K-means. However, selecting redundant low-confidence samples may harm the effectiveness of model validation. This motivates us to select samples approximating the testing data distribution, similar to the objective of active testing (Kossen et al., 2021). Biasing towards either criterion would result in a suboptimal HILTTA. Our design mainly aims to balance the preference over selecting low-confidence and diverse samples.
A similar objective as above was achieved by a recent active learning approach, BADGE (Ash et al., 2020), which takes the gradient of loss w.r.t. the penultimate layer feature as the uncertainty weighted feature as $g_{x}=\frac{\partial}{\partial\theta_{\theta_{out}}}\mathcal{L}_{ce}(h(x;\theta),\hat{y})$, where $\theta_{out}\in\mathbbm{R}^{D\times C}$ is the classifier weights. Hence, the gradient embedding $g_{x}\in\mathbbm{R}^{D\cdot C}$ is of dimension $D\cdot C$. When the semantic space is substantially high, e.g. $C=1000$ for ImageNet (Deng et al., 2009), the above gradient embedding results in extremely high dimension features. For instance, a typical penultimate layer feature is 2048 channels and the gradient embedding could reach 2,048,000 channels. Such a high dimension prevents efficient clustering operations, e.g. due to memory constraints, for sampling. To mitigate the ultra-high dimension challenge, we integrate uncertainty into feature representation by multiplying the uncertainty, measured by the margin confidence, to the feature representation as in Eq. 7, where $\mathrm{sort}^{d}$ refers to the descending sorting.
@@
The above design prioritizes the difference between the probability of the most confident and second most confident classes as a quantification of model uncertainty. Such a design is superior to alternative designs, e.g. entropy, for the following reasons. First, deep learning models tend to be over confident with low entropy in general. Thus the gap between top 1 and top 2 class probabilities may better suggest the model confidence than entropy. Moreover, the entropy may become less distinct for a classification task with many categories as the randomness on non-ground-truth classes may affect the overall entropy calculation. Finally, our empirical observations in Sect. 4.3 also suggest the superiority of using the probability margin.
@@
Our K-Margin approach with the K-Center clustering approach offers a computationally efficient alternative to K-Means++ while similarly balancing uncertainty and diversity in sample selection. Unlike K-Means++, which requires iterative center updates, K-Center clustering selects points that maximize coverage by minimizing the maximum distance from each sample to its nearest center. This method provides a representative subset with less computational overhead, which is more suitable in TTA settings. By focusing on high-uncertainty, diverse samples, K-Center clustering ensures a labeled subset that accurately reflects the test distribution. It guarantees that the labeled subset spans the feature space with fewer computations, thereby maintaining a consistent approximation of the test distribution. This enhances model selection and supervised training by creating a robust validation set that captures broad feature variance without redundant low-confidence samples.
@@ 4.1 Experimental setting
We use the optimizer recommended in the original papers for each method’s unsupervised training. We set the momentum $\beta$ for 0.5. For supervised training, we use the Adam optimizer with a learning rate of 1e-5. We consider seven distinct values for each hyper-parameter category regarding model selection, detailed in Tab. 8.
@@ 4.2 Evaluations on HILTTA
Table 2: Average classification error (lower is better) through ongoing adaptation to varying domains. "Worst", "Avg.", and "Best" indicate the worst, average, and best performance within the candidate hyper-parameter pool $\Omega$. The improvement of error rate for each off-the-shelf TTA method after combining with HIL is appended in the bracket as ($\delta$). * denotes methods under manually tuned hyper-parameters. $\diamondsuit$ denotes methods combined with our proposed HILTTA (with model selection). We adapt VeSSAL for ATTA by combining TENT as an unsupervised training method.
@@ 4.3 Ablation & Additional Study
Evaluations beyond 2D Image Recognition: To further evaluate the effectiveness of our method beyond 2D image classification task, we conduct experiments on 3D point cloud classification using DGCNN (Wang et al., 2019) as the backbone, adapting it continuously across 15 domains within the ModelNet40-C (Sun et al., 2022) dataset. We maintained the same hyper-parameter candidate set as outlined in Tab. 8 and followed the experimental setup detailed in Tab. 2, with a batch size of 32 and an annotation percentage of 3.2%. The results, presented in Fig. 3, show that unsupervised TTA still consistently outperforms the baseline without any adaptation (Source). However, due to the sensitivity to hyper-parameters, all unsupervised TTA methods exhibit a substantial variation in performance with the “Best” and “Worst” hyper-parameters. All off-the-shelf TTA methods are consistently improved when integrated with our HILTTA method. Five out of the six methods witness a more significant boost in performance, surpassing their unsupervised counterpart with “Best” hyper-parameters.
@@
Performance Under Adjustable Human Intervention Frequency: We investigate the effectiveness of HILTTA under varying human intervention frequencies (N), as shown in Tab. 7. Instead of requesting human intervention for every batch of test streaming data, we reduce the frequency by only requesting intervention for every N-th batch. Consequently, the label rate is reduced to 1/N. Model selection is also performed in every N batch, while for batches without human intervention, the previously selected hyperparameters are used for test-time adaptation. Remarkably, even when $N=10$, corresponding to a mere 0.3% label rate, HILTTA combined with TENT significantly outperforms the original TENT without human intervention, achieving an error rate of 18.61% compared to 27.77%. This demonstrates that even with a very limited labeling budget and sparse human intervention, notable performance gains can be achieved. By employing an adjustable human intervention frequency, there is no need to annotate every batch, further enhancing HILTTA’s practical impact.
Table 7: Performance of TENT (Wang et al., 2020) with HILTTA under different human intervention frequency (N) on CIFAR10-C.
Human Intervention Frequency (N) W/O HIL 1 2 3 4 5 6 7 8 9 10
Error Rate (%) 27.77 15.87 16.43 17.15 17.53 17.94 17.99 18.00 18.05 18.39 18.61
@@ Appendix B Additional Experimental Analysis
B.1 Ablation Study on the Impact of $\beta$
We evaluate the impact of the moving average momentum parameter, $\beta$, in Eq. 6 when applying TENT (Wang et al., 2020) combined with our proposed HILTTA. The experiments were conducted across four datasets: ImageNet-C, ImageNet-D, CIFAR100-C, and CIFAR10-C. We varied $\beta$ over a wide range, from 0.1 to 0.9, and observed in Fig. 7 that the performance remained robust throughout. This demonstrates that our proposed HILTTA method is not sensitive to the choice of $\beta$ values, maintaining stability across different settings.
Figure 7: Performance of TENT (Wang et al., 2020) combined with our proposed HILTTA under different $\beta$ on 4 different datasets.
B.2 Comparison to Bi-level Optimization
In some scenarios, the cost of collecting annotated data can outweigh the computational expense, making hyperparameter optimization via bi-level optimization (Liu et al., 2018) a viable approach. However, bi-level optimization has limitations, particularly its sensitivity to the meta-learning rate (the learning rate in the outer loop). We compare our method, HILTTA, with the approach proposed in (Liu et al., 2018), using Adam as the meta-optimizer. The Higher package (Grefenstette et al., 2019) is adopted for meta optimization. As shown in Fig 8, while the best performance achieved by bi-level optimization is comparable to ours (59.34% vs. 58.35% error rate), bi-level methods can be prone to collapse if the meta-learning rate is not carefully tuned. Additionally, bi-level optimization is unable to optimize hyperparameters that are non-differentiable.
Figure 8: Performance of TENT (Wang et al., 2020) combined with our proposed HILTTA and bi-level optimization on the ImageNet-C dataset.
B.3 Comparison to Semi-supervised Learning Methods
To evaluate the effectiveness of our proposed HILTTA, we compare it against semi-supervised methods, such as FixMatch (Sohn et al., 2020) and MeanTeacher (Tarvainen & Valpola, 2017), both using the same label rate of 3%. Additionally, we include ATTA methods like SimATTA (Gui et al., 2024) and VeSSAL (Saran et al., 2023), adapted to the ATTA setting by integrating them with TENT (Wang et al., 2020), also with label rate of 3%. As shown in Tab. 9, our HILTTA (HIL+TENT) surpasses all other TTA, semi-supervised learning, and ATTA approaches. Semi-supervised methods typically require longer training times and multiple epochs to achieve convergence, making them less effective in stream-based tasks. It is worth noting that while both semi-supervised and ATTA methods rely on manually tuned hyperparameters, HILTTA achieves superior performance with model selection.
Table 9: HILTTA Comparison with Semi-supervised Learning and Active TTA methods.
Setting TTA Semi-supervised Learning ATTA HILTTA(Ours)
Method TENT FixMatch MeanTeacher SimATTA VeSSAL HIL + TENT
Error Rate(%) 27.77 17.03 16.59 17.66 17.47 15.87
Although semi-supervised learning and ATTA methods can achieve performance comparable to HILTTA, they suffer from significant sensitivity to hyperparameters. As shown in Fig. 9, we compare the performance of TENT, HIL+TENT, and FixMatch on the CIFAR10-C dataset with learning rate as the variable hyperparameter. The results indicate that semi-supervised methods such as FixMatch are particularly sensitive to changes in learning rate, highlighting the challenges of tuning these methods effectively. Our proposed HILTTA is not only aimed at improving performance but, more importantly, at addressing hyperparameter sensitivity issues by synergizing active learning with model selection.
Figure 9: Performance of Fixmatch (Sohn et al., 2020) and TENT (Wang et al., 2020) combined with our proposed HILTTA on the CIFAR10-C dataset.
B.4 Studies on the Role of Regularization in HILTTA
We evaluate the impact of the proposed anchor regularization and EMA smoothing techniques for model selection within the HILTTA framework. These methods are designed to prevent the model from overfitting to the local data distribution in the testing data stream. While removing these regularizations may improve adaptation to a single domain, it negatively impacts performance across future domains. As shown in Tab. 10, in the continual TTA setting, model selection without regularization results in significantly worse performance compared to when regularization is applied (17.25% vs. 15.87%). This indicates that regularization is crucial for maintaining robust performance in continual TTA.
@@ B.4 Studies on the Role of Regularization in HILTTA
B.5 Comparing with Knowledge Distillation and Ensemble Learning
We carried out additional experiments by comparing with ensemble learning and multi-teacher knowledge distillation methods. Specifically, for ensemble learning, we average the posterior predicted by each candidate model and make a prediction with the average posterior. For multi-teacher knowledge distillation, we use the pseudo label predicted by the ensemble model for self-training. As observed from Tab. 11, both ensemble learning and multi-teacher knowledge distillation yields inferior results than HILTTA. This is most likely caused by incorporating poor candidate models (ensemble learning) and using poor pseudo labels for self-training, i.e. learning from noisy labeled data (multi-teacher knowledge distillation).
@@ B.5 Comparing with Knowledge Distillation and Ensemble Learning
B.6 Comparing Adaptation Time & Annotation Time
We carried out empirical studies by measuring the wall-clock time lapses of the inference step (Inference Time), adaptation step (Adaptation Time), and human annotation (Annotation Time) per sample. In specific, the "Inference Time" refers to the time-lapse of pure forward pass, the "Adaption Time" refers to the time-lapse of backpropagation and weights update, and "Annotation Time" refers to the estimated time lapse required for human to annotate images at a 3% labeling rate. We evaluate under three categories of annotation cost, i.e. 2 sec/sample, 5 sec/sample, and 10 sec/sample. As seen from Tab. 12, despite the adaptation time being higher than the inference time (8.7 ms v.s. 0.4 ms), the annotation time is the most expensive step. These results suggest model selection and adaptation do not slow down the throughput under the HILTTA protocol.